Merry had $20,000 to invest. she invested part of this money in bonds paying 10% annual simple interest and the rest of the money in a savings account giving 5%annual interest. at the end of the year,she received $1,800as extra income. how much money did merry place in the savings account?
Answer by mananth(16946)

Part I 10.00% per annum ------------- Amount invested =x
Part II 5.00% per annum ------------ Amount invested = y
20000
Interest----- 1800.00

Part I 10.00% per annum ---x
Part II 5.00% per annum ---y
Total investment
x + 1 y= 20000 -------------1
Interest on both investments
10.00% x + 5.00% y= 1800
Multiply by 100
10 x + 5 y= 180000.00 --------2
Multiply (1) by -10
we get
-10 x -10 y= -200000.00
Add this to (2)
0 x -5 y= -20000
divide by -5
y = 4000
Part I 10.00% $ 16000
Part II 5.00% $ 4000

CHECK
16000 --------- 10.00% ------- 1600.00
4000 ------------- 5.00% ------- 200.00
Total -------------------- 1800.00